OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (anonymous):

First find your slope using the formula \[\frac{ y2 -y1 }{ x2 -x1 }\] after that choose any of the two co-ordinates and plug it into this formula. \[(y -yo) =m (x-xo)\] Where yo is your y -coordinate and xo is your x-coordinate and m is the slope. solve for y =mx +b...

OpenStudy (anonymous):

can you type answer please

OpenStudy (anonymous):

haha ok I'll show you how to do it.

OpenStudy (anonymous):

because I'm bad at those i know how to do it but it will take me 10 min to solve one

OpenStudy (anonymous):

hey! that no excuse I was the same at the beginning as well .... just takes practice.

OpenStudy (anonymous):

Math is not really favorite subject to me its on second place

OpenStudy (anonymous):

LOL I think math sucks but its interesting haha...

OpenStudy (anonymous):

Can help me solve its last question for the assignment

OpenStudy (anonymous):

2x -y -3 = 0

OpenStudy (anonymous):

thanks :) i have problems with negatives most of them i know

OpenStudy (anonymous):

yw keep practicing alright...